OpenStudy (anonymous):

Convert V=x^3 to x=...?

OpenStudy (jhannybean):

solving for x instead?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

You want to isolate the x, to "undo" the cube, you have to take the cube root of both sides. Can you tell me what you can?

OpenStudy (jhannybean):

what you get*

OpenStudy (anonymous):

What would taking the cube of both sides look like though?

OpenStudy (jhannybean):

\[\large \sqrt[3]{V} = \sqrt[3]{x^3}\]

OpenStudy (jhannybean):

When you take the cube root of x, it turns into : \[(x^3)^{1/3} = x^{3 \cdot 1/3}\]

OpenStudy (jhannybean):

and what is \[3 \cdot \frac{1}{3}\]

OpenStudy (anonymous):

1

OpenStudy (jhannybean):

Good :)

OpenStudy (jhannybean):

So what would be your answer?

OpenStudy (anonymous):

I'm honestly still confused. Would it be just the 3 square root V?

OpenStudy (jhannybean):

Yes it would :)

OpenStudy (jhannybean):

What's confusing you?

OpenStudy (jhannybean):

Basically, if we are given an equation ans we are solving for one variable in place of the other, we want to "isolate" the variable we are solving for.

OpenStudy (jhannybean):

In this case, x was raised to the power of 3, so we had to "undo" its power to get it all by itself.

OpenStudy (jhannybean):

and the only way to "undo" \(x^3\) is to take it's inverse power, which is \(\large \sqrt[3]{x}\)

OpenStudy (anonymous):

Ohhh ok, I was confused about the x part, but now I get it... I have one more problem, can you help me with it?

OpenStudy (jhannybean):

And that is why you got x o the power of 1, because \(3 \cdot \frac{1}{3} = 1\)

OpenStudy (jhannybean):

Sure thing :)

OpenStudy (anonymous):

Thank you! So the square root of a^2+b^2=c

OpenStudy (anonymous):

And we are solving for a

OpenStudy (jhannybean):

Ok, you want to isolate a^2 on the left side of the equation first, how would you do that?

OpenStudy (anonymous):

Wouldn't you square both sides first to get rid of the square root?

OpenStudy (jhannybean):

Nope, that would make the left side more complicated to solve than need be, try again :P

OpenStudy (anonymous):

Subtract b^2?

OpenStudy (jhannybean):

Good :)

OpenStudy (jhannybean):

And what do you get?

OpenStudy (anonymous):

a^2=c^2-b^2

OpenStudy (jhannybean):

Good:)

OpenStudy (jhannybean):

Now you can take the square root of both sides, just like your previous problem, "undo" the square on a^2.

OpenStudy (anonymous):

Wait, was I supposed to have c^2?

OpenStudy (jhannybean):

oh no, you don't need to square that. You just leave the c term as i is.

OpenStudy (jhannybean):

it*

OpenStudy (anonymous):

So would it be a=square root of c-b^2, or would you take the b^2 out?

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

You don't need tot take the b out.